Use linear substitution to solve the following first-order differential equation


𝑑𝑦/𝑑π‘₯=(2π‘₯+𝑦)/(2π‘₯+𝑦+1)

1
Expert's answer
2021-11-02T15:32:59-0400

"\frac{d y}{d x}=\frac{2 x+y}{2 x+y+1} \ ...(1)
\\\text{By substitution }
\\\text{let}\ v=2 x+y
\\\text { then } 
\frac{d v}{d x} =2+\frac{d y}{d x} \\
\text { so } \frac{d y}{d x} =\frac{d v}{d x}-2
\\\therefore \text{From (1) we get}
\\\frac{d v}{d x}-2=\frac{v}{v+2} 
\\\frac{d v}{d x}=\frac{v}{(v+1)}+2 
=\frac{v+2(v+1)}{(v+1)} \\
\frac{d v}{d x}=\frac{3 v+2}{(v+1)}

\\\therefore \quad \frac{(v+1)}{(3 v+2)} d v=d x
\\\text{Integrating both sides}
\\\int \frac{(v+1)}{(3 v+2)} d v=\int d x+c
\\\frac{1}{3} \int\left[\frac{3 v+2+1}{3 v+2}\right] d v=x+c
\\\Rightarrow \frac{1}{3} \int\left[1+\frac{1}{(3 v+2)}\right] d v=x+c
\\\Rightarrow \quad \int 1 d v+\frac{1}{3} \int \frac{d v}{\left(v+\frac{2}{3}\right)}=3 x+c
\\\Rightarrow \quad v+\frac{1}{3} \ln|v+\frac{2}{3}|=3 x+c
\\\text{Putting}\ v=2 x+y\ \text{we have}
\\2 x+y+\frac{1}{3} \ln \left|2 x+y+\frac{2}{3}\right|=3 x+c \\

\\\Rightarrow \quad y+\frac{1}{3} \ln \left|2 x+y+\frac{2}{3}\right|= x+C \\
\\\Rightarrow \quad y+\frac{1}{3} \ln \left|2 x+y+\frac{2}{3}\right|-x=C"